# RouteKG: A knowledge graph-based framework for route prediction on road networks

### v2

@@ 4.3.3 Future Route Prediction through Knowledge Graph Completion
4.4 Route Generation Module
Given predicted future route probabilities $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}})$, the Route Generation Module generates multiple possible future routes from these probabilities. An $n$-ary tree-based algorithm, Spanning Route, is proposed to generate these future routes based on the predicted probabilities. This algorithm is visualized in Figure 5 through a simplified case where $n=2$ and only $\Gamma=3$ predicted future links are illustrated. For each tree node, we designate four attributes: name, parent, end_node, and pred. The name corresponds to the identification of the leaf, while the parent points to the predecessor of the current leaf. The attribute end_node signifies the terminal node of the current predicted link, and pred embodies the present predictions.
Input :  $\Gamma^{\prime}$ probability distributions $\left\{\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})\right\}_{\gamma=1}^{\Gamma^{\prime}}$;
road network $\mathbf{G}=(\mathbf{V},\mathbf{E})$;
NAE matrix $\mathbf{A}$;
the tree’s degree $n$.
Output :  Top-$K$ predicted future routes $\left\{\widetilde{x_{i,k}^{f}}\right\}_{k=1}^{K}$.
1 root $\leftarrow$ CreateNewNode(name = “root”, parent = NIL, end_node = $v_{i,\Gamma}^{s}$, pred = NIL)
@@ 16       $\text{path}_{k}$ = Traverse(root, leaves[$k$])
To formally introduce the Spanning Route algorithm, we provide the pseudo-code for generating the multiple future routes based on the predicted probabilities in Algorithm 1. Specifically, the algorithm encompasses four primary functions. The CreateNewNode function instantiates a new node in the tree given its attributes, while the GetLeaves function takes the root node as input and outputs all leaves of the tree. The GetTopK function retrieves the top-$k$ predictions given a predicted probability and $k$, and the Traverse function applies a tree-based Depth-First Search (DFS) traversal algorithm—specifically a Pre-Order traversal (Tarjan 1972)—to acquire the path from the root to a specified leaf. This last function is instrumental in merging the predicted links into cohesive predicted routes. We note that certain detailed masking and indexing operations have been omitted in the presented pseudo code for clarity. For more details, please refer to the minibatch version of the Spanning Route algorithm 3 detailed in the appendix (B).
4.5 Rank Refinement Module
The top-$K$ future routes candidates $\widetilde{\mathcal{X}^{f}}=\left\{\widetilde{\mathcal{X}^{f}_{k}}\right\}_{k=1}^{K}=\left\{\left\{\widetilde{x_{i,k}^{f}}\right\}_{k=1}^{K}\right\}_{i=1}^{|\mathcal{X}^{f}|}$ offer an initial selection of possible future routes. However, the dependencies among different links within these routes are solely based on the connectivity of the road network. Given that the consistency and other spatial relations of links within a route also affect people’s choices of routes, a more refined approach is needed for accurate future route prediction. To achieve this, we leverage learned spatial relations $\mathcal{R}$ to model the dependencies between different links and rerank the candidate routes based on the learned dependencies. This process can be denoted as $\left\{\widehat{\mathcal{X}^{f}_{k}}\right\}_{k=1}^{K}=\mathcal{M}_{r}\left(\left\{\widetilde{\mathcal{X}^{f}_{k}}\right\}_{k=1}^{K},\mathcal{R};\Theta_{r}\right)$
@@ 4.5 Rank Refinement Module
$\mathbf{r}^{f,c}_{i,margin}$ $=\frac{1}{\Gamma^{\prime}-1}\sum_{j=1}^{\Gamma^{\prime}-1}\widetilde{\mathbf{x}^{f}_{i,\perp^{c}}}\left[:,j,:\right]-\widetilde{\mathbf{x}^{f}_{i,\perp^{c}}}\left[:,j+1,:\right]$ (12)
@@
$\mathbf{r}^{f,s}_{i,margin}$ $=\frac{1}{\Gamma^{\prime}-1}\sum_{j=1}^{\Gamma^{\prime}-1}\widetilde{\mathbf{x}^{f}_{i,\perp^{s}}}\left[:,j,:\right]-\widetilde{\mathbf{x}^{f}_{i,\perp^{s}}}\left[:,j+1,:\right],$
@@
where $\mathbf{r}^{f,c}_{i,margin}\in\mathbb{R}^{K\times\delta_{\mathcal{E}}}$ is the connectivity margin and $\mathbf{r}^{f,s}_{i,margin}\in\mathbb{R}^{K\times\delta_{\mathcal{E}}}$ the consistency margin. Following this, the derived $\mathbf{r}^{f,c}_{margin}$ and $\mathbf{r}^{f,s}_{margin}$ are flattened and, together with the flattened route embedding $\widetilde{\mathbf{x}^{f}_{i}}\in\mathbb{R}^{K\cdot\Gamma^{\prime}\cdot\delta_{\mathcal{E}}}$ and route direction embedding $\widetilde{\mathbf{x}^{f,d}_{i}}\in\mathbb{R}^{K\cdot\Gamma^{\prime}\cdot\delta_{\mathcal{E}}}$, used to compute the new rank:
$\mathrm{Pr}\left(\widetilde{R}\right)=\mathrm{Softmax}\left(\mathrm{MLP}_{r}\left(\mathbf{r}^{c}_{margin}\text{ }\|\text{ }\mathbf{r}^{s}_{margin}\text{ }\|\text{ }\mathrm{MLP}_{f}\left(\widetilde{\mathbf{x}^{f}_{i}}\text{ }\|\text{}\widetilde{\mathbf{x}^{f,d}_{i}}\right)\right)\right),$ (13)
@@
where $\mathrm{Pr}(\widetilde{R})\in\mathbb{R}^{K}$ denotes the probability distribution over the $K$ predicted future routes being the actual future routes. Based on this probability, we can determine the new predicted rank, resulting in reranked future route predictions denoted as $\left\{\widehat{x^{f}_{i,k}}\right\}_{k=1}^{K}$. We also adopt the cross-entropy loss for the Rank Refinement Module:
$\mathcal{L}_{rank}=-\log\mathrm{Pr}\left(\widetilde{R}\right)\left[x^{f}_{i}\right],$ (14)
@@
Note that samples from the set of multiple predicted future routes are excluded if they do not contain a ground truth future route. While our KG facilitates an effective overall selection (i.e., top-$K$) of future routes, it is crucial to notice that an enhancement in the top-$K$ predictions does not necessarily translate to a superior top-$1$ prediction. Therefore, we directly refine the top-$1$ prediction using the initial predictions in our implementation. This is done by employing a MLP to encode the initial prediction embeddings $\mathbf{x}^{f}_{i}$ and $\mathbf{x}^{f,d}_{i}$, along with the last observed link $\mathbf{e}^{\Gamma}_{i}$ and $\mathbf{e}^{d,\Gamma}_{i}$, and estimated goal direction $\mathbf{r}^{d}_{i}$:
$\widetilde{x^{f}_{i}}=\mathrm{MLP}_{k}\left(\mathbf{e}^{\Gamma}_{i}\text{ }\|\text{ }\mathbf{e}^{d,\Gamma}_{i}\text{ }\|\text{ }\mathbf{r}^{d}_{i}\text{ }\|\text{ }\mathrm{MLP}_{x}\left(\mathbf{x}^{f}_{i}\text{ }\|\text{ }\mathbf{x}^{f,d}_{i}\right)\right),$ (15)
@@
where $\widetilde{x^{f}_{i}}\in\mathbb{R}^{\Gamma^{\prime}\times|\mathcal{E}|}$, is also optimized by minimizing the corresponding cross entropy loss. Subsequently, the top-$1$ prediction is generated using the Route Generation Module for the $n=1$ case. The generated future route can be inserted into the first position, replacing the original $K$-th prediction, to obtain the refined top-$K$ future route predictions $\left\{\widehat{x^{f}_{i,k}}\right\}_{k=1}^{K}$.
4.6 Multi-Objectives Optimization
@@
where $w_{\cdot}$ are weights for different loss.
The complete learning procedure of RouteKG are detailed in Algorithm 2.
@@
1 Randomly initialize $\Theta_{kg}=\left\{\mathbf{W}_{\mathcal{E}},\mathbf{W}_{\mathcal{R}^{c}},\mathbf{W}_{\mathcal{R}^{s}},\mathbf{W}_{\mathcal{R}^{a}},\mathbf{W}_{\mathcal{R}^{d}},\mathbf{P}_{\mathcal{R}^{c}},\mathbf{P}_{\mathcal{R}^{s}},\mathbf{P}_{\mathcal{R}^{a}},\mathbf{P}_{\mathcal{R}^{d}},\mathrm{MLP}_{d}\right\}$ and $\Theta_{r}=\left\{\mathrm{MLP}_{r},\mathrm{MLP}_{f},\mathrm{MLP}_{k},\mathrm{MLP}_{x}\right\}$.
@@
7       $\left\{\widetilde{\mathcal{X}^{f}_{k}}\right\}_{k=1}^{K}\leftarrow\mathcal{M}_{g}\left(\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathbf{G},\mathbf{A}\right)$
8       $\left\{\widehat{\mathcal{X}^{f}_{k}}\right\}_{k=1}^{K}\leftarrow\mathcal{M}_{r}\left(\left\{\widetilde{\mathcal{X}^{f}_{k}}\right\}_{k=1}^{K},\mathcal{R};\Theta_{r}\right)$
@@ 5.1 Data
The Shanghai dataset consists of trajectory records from 10,609 taxis from April 16, 2015, to April 21, 2015, with an average sampling rate of approximately 10 seconds per record. We concentrated on a specific region in Shanghai, adhering to the parameters outlined by (Zhao & Liang 2023). The chosen region’s road network incorporates 320 nodes and 714 links. Each data entry includes the taxi ID, date, time, longitude, latitude, and an occupied flag indicator.
For data preprocessing, we initially employed a fast map-matching algorithm (Yang & Gidofalvi 2018) to convert GPS traces into routes on the respective road network. We then cleaned the data, eliminating routes that contained loops and those with too few links (i.e., less than ten links). Subsequently, the refined Chengdu dataset contained 93,125 routes, while the Shanghai dataset comprised 24,468 routes. The road networks of Chengdu and Shanghai are visually represented in Figure 6, and the key network statistics are summarized in Table 2.
Figure 6: The selected road networks in Chengdu and Shanghai.
Table 2: Summary statistics of road networks in Chengdu and Shanghai, where ID & OD refer to in-degree and out-degree, respectively.
@@
5.2 Baseline Methods
In this study, we compare our approach with several established baselines to evaluate performance. These baselines include:
@@ 5.2 Baseline Methods
  • Dijkstra (Dijkstra 1959): Dijkstra’s algorithm is a prominent method for finding the shortest paths between nodes in a graph, where the path length is assumed to be the sum of link length to reflect realistic geographic distance. This baseline can only work when the exact goal location is given.
  • RNN (Rumelhart et al. 1986): The RNN is an artificial neural network that recognizes patterns in sequential data. It accomplishes this by utilizing internal memory to process arbitrary sequences of inputs, making it effective for predicting future routes.
  • GRU (Cho et al. 2014): The GRU is a type of RNNs that utilizes gating mechanisms to capture long-term dependencies in the data, thereby improving the model’s predictive capabilities. This is especially beneficial for applications in trajectory prediction, where long-term dependencies play a crucial role.
  • LSTM (Hochreiter & Schmidhuber 1997): The LSTM is another variant of RNNs that addresses the problem of learning long-term dependencies in data. It does this by implementing a special architecture consisting of a series of memory cells, which effectively control the flow of information, making LSTMs advantageous for predicting future routes.
  • NetTraj (Liang & Zhao 2021): NetTraj is an advanced network-based trajectory prediction model specifically designed for predicting future movements in road networks. It leverages the inherent structure of road networks and utilizes historical trajectory data for accurate predictions. By integrating the Graph Attention Network (GAT) with LSTM, NetTraj offers a robust framework for making future trajectory predictions.
  • RCM-BC (Zhao & Liang 2023): The Route Choice Model-Behavioral Cloning (RCM-BC) is a behavioral cloning approach designed for route choice modeling in sequential decision-making scenarios. It employs supervised learning to create a policy that maps states to actions based on observed behavior to predict future routes. This baseline also requires the knowledge of the exact goal location to work.
Table 3: Performance comparison of different methods.
@@
5.3 Main Results
5.3.1 Experimental Settings
To comprehensively assess model performance, we design experiments based on the three subproblems as defined in Section 3.1: (1) route prediction with unknown goal $\mathcal{F}_{1}$, (2) route prediction with goal direction only $\mathcal{F}_{2}$, and (3) route prediction with complete goal information $\mathcal{F}_{3}$. We refer to these three subproblems as NoGoal, GoalD, and Goal. They reflect varying degrees of information availability regarding the road user’s intended destination, and represent a broad range of real-world application scenarios. For instance, a system might not know the user’s exact destination due to privacy concerns but could have access to more general information, such as the goal direction.
@@ 5.3.1 Experimental Settings
We consider the top-$k$ predictions for the $i$-th observed route $\left\{\widehat{x^{f}_{i,k}}\right\}_{k=1}^{K}$=$\left\{\left\{\widehat{e_{i,k}^{j}}\right\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\right\}_{k=1}^{K}$ and the actual $i$-th future route $x_{i}^{f}=\left\{e_{i}^{j}\right\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$ to define our evaluation metrics. The link-level recall R@K is defined as
$R@K=\frac{1}{|\mathcal{X}|}\sum_{i=1}^{|\mathcal{X}|}\max_{k=1}^{K}\left[\frac{1}{\Gamma^{\prime}}\sum_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\mathbb{I}\left(\widehat{e_{i,k}^{j}}=e_{i}^{j}\right)\right].$ (17)
where $\mathbb{I}(\cdot)$ is the indicator function: $\mathbb{I}(a=b)=\begin{cases}1&\text{if }a=b\\ 0&\text{otherwise}\end{cases}$.
Similarly, the route-level recall R@K is defined as
$R@K=\frac{1}{|\mathcal{X}|}\sum_{i=1}^{|\mathcal{X}|}\max_{k=1}^{K}\left[\mathbb{I}\left(\sum_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\mathbb{I}\left(\widehat{e_{i,k}^{j}}=e_{i}^{j}\right),\Gamma^{\prime}\right)\right].$ (18)
We also compute the route-level MRR of the top-$k$ predictions, M@K, as follows:
$M@K=\frac{1}{|\mathcal{X}|}\sum_{i=1}^{|\mathcal{X}|}\sum_{k=1}^{K}\frac{1}{k}\left[\mathbb{I}\left(\sum_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\mathbb{I}\left(\widehat{e_{i,k}^{j}}=e_{i}^{j}\right),\Gamma^{\prime}\right)\right].$ (19)
The experiments are conducted on an Ubuntu server leveraging the Python 3.6 environment. The deep learning computations are performed using the PyTorch framework. The server’s hardware specifications include an Intel(R) Xeon(R) Platinum 8375C CPU with a clock speed of 2.90GHz, coupled with 8 NVIDIA GeForce RTX 3090 GPUs, each featuring 24GB of memory. To ensure the robustness and generalizability of our model, hyperparameters are tuned based on the performance of the validation set. The fine-tuning of these hyperparameters is crucial for balancing the bias-variance trade-off and optimizing the model’s performance. All hyperparameters are listed in C.
@@
Figure 7: Example results in Chengdu (upper) and Shanghai (lower), with the red line indicating last observed link, green line the actual future route, blue lines the predicted future routes.
5.3.2 Main Results Analysis

### v3

@@ IV-D Route Generation Module
1
Input :  $\Gamma^{\prime}$ probability distributions $\{\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})\}_{\gamma=1}^{\Gamma^{\prime}}$; road network $\mathbf{G}=(\mathbf{V},\mathbf{E})$;
NAE matrix $\mathbf{A}$; the tree’s degree $n$.
Output :  Top-$K$ predicted future routes $\{\widetilde{x_{i,k}^{f}}\}_{k=1}^{K}$.
2
3root $\leftarrow$ CreateNewNode(name = “root”, parent = NIL, end_node = $v_{i,\Gamma}^{s}$, pred = NIL)
@@ 16       $\text{path}_{k}$ = Traverse(root, leaves[$k$])
To formally introduce the Spanning Route algorithm, we provide the pseudo-code for generating the multiple future routes based on the predicted probabilities in Algorithm 1. Specifically, the algorithm encompasses four primary functions. The CreateNewNode function instantiates a new node in the tree given its attributes, while the GetLeaves function takes the root node as input and outputs all leaves of the tree. The GetTopK function retrieves the top-$k$ predictions given a predicted probability and $k$, and the Traverse function applies a tree-based Depth-First Search (DFS) traversal algorithm, specifically a Pre-Order traversal [76]—to acquire the path from the root to a specified leaf. This last function is instrumental in merging the predicted links into cohesive predicted routes. We note that certain detailed masking and indexing operations have been omitted in the presented pseudo code for clarity. For more details, please refer to the minibatch version of the Spanning Route algorithm 3 detailed in the appendix (VI-A).
IV-E Rank Refinement Module
The top-$K$ future routes candidates $\widetilde{\mathcal{X}^{f}}=\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}=\{\{\widetilde{x_{i,k}^{f}}\}_{k=1}^{K}\}_{i=1}^{|\mathcal{X}^{f}|}$ offer an initial selection of possible future routes. However, the dependencies among different links within these routes are solely based on the connectivity of the road network. Given that the consistency and other spatial relations of links within a route also affect people’s choices of routes, a more refined approach is needed for accurate future route prediction. To achieve this, we leverage learned spatial relations $\mathcal{R}$ to model the dependencies between different links and rerank the candidate routes based on the learned dependencies. This process can be denoted as $\{\widehat{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}=\mathcal{M}_{r}(\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K},\mathcal{R};\Theta_{r})$
@@ IV-E Rank Refinement Module
$\mathbf{r}^{f,c}_{i,m}$ $=\frac{1}{\Gamma^{\prime}-1}\sum_{j=1}^{\Gamma^{\prime}-1}\widetilde{\mathbf{x}^{f}_{i,\perp^{c}}}\left[:,j,:\right]-\widetilde{\mathbf{x}^{f}_{i,\perp^{c}}}\left[:,j+1,:\right]$ (12)
@@
$\mathbf{r}^{f,s}_{i,m}$ $=\frac{1}{\Gamma^{\prime}-1}\sum_{j=1}^{\Gamma^{\prime}-1}\widetilde{\mathbf{x}^{f}_{i,\perp^{s}}}\left[:,j,:\right]-\widetilde{\mathbf{x}^{f}_{i,\perp^{s}}}\left[:,j+1,:\right],$
@@
where $\mathbf{r}^{f,c}_{i,m}\in\mathbb{R}^{K\times\delta_{\mathcal{E}}}$ is the connectivity margin and $\mathbf{r}^{f,s}_{i,m}\in\mathbb{R}^{K\times\delta_{\mathcal{E}}}$ the consistency margin. Following this, the derived $\mathbf{r}^{f,c}_{m}$ and $\mathbf{r}^{f,s}_{m}$ are flattened and, together with the flattened route embedding $\widetilde{\mathbf{x}^{f}_{i}}\in\mathbb{R}^{K\cdot\Gamma^{\prime}\cdot\delta_{\mathcal{E}}}$ and route direction embedding $\widetilde{\mathbf{x}^{f,d}_{i}}\in\mathbb{R}^{K\cdot\Gamma^{\prime}\cdot\delta_{\mathcal{E}}}$, used to compute the new rank:
$\mathrm{Pr}(\widetilde{R})=\mathrm{Softmax}(\mathrm{MLP}_{r}(\mathbf{r}^{c}_{m}\text{ }\|\text{ }\mathbf{r}^{s}_{m}\text{ }\|\text{ }\mathrm{MLP}_{f}(\widetilde{\mathbf{x}^{f}_{i}}\text{ }\|\text{ }\widetilde{\mathbf{x}^{f,d}_{i}}))),$ (13)
@@
where $\mathrm{Pr}(\widetilde{R})\in\mathbb{R}^{K}$ denotes the probability distribution over the $K$ predicted future routes being the actual future routes. Based on this probability, we can determine the new predicted rank, resulting in reranked future route predictions denoted as $\{\widehat{x^{f}_{i,k}}\}_{k=1}^{K}$. We also adopt the cross-entropy loss for the Rank Refinement Module:
$\mathcal{L}_{rank}=-\log\mathrm{Pr}(\widetilde{R})\left[x^{f}_{i}\right],$ (14)
@@
Note that samples from the set of multiple predicted future routes are excluded if they do not contain a ground truth future route. While our KG facilitates an effective overall selection (i.e., top-$K$) of future routes, it is crucial to notice that an enhancement in the top-$K$ predictions does not necessarily translate to a superior top-$1$ prediction. Therefore, we directly refine the top-$1$ prediction using the initial predictions in our implementation. This is done by employing an MLP to encode the initial prediction embeddings $\mathbf{x}^{f}_{i}$ and $\mathbf{x}^{f,d}_{i}$, the last observed link $\mathbf{e}^{\Gamma}_{i}$ and $\mathbf{e}^{d,\Gamma}_{i}$, and estimated goal direction $\mathbf{r}^{d}_{i}$:
$\widetilde{x^{f}_{i}}=\mathrm{MLP}_{k}(\mathbf{e}^{\Gamma}_{i}\text{ }\|\text{}\mathbf{e}^{d,\Gamma}_{i}\text{ }\|\text{ }\mathbf{r}^{d}_{i}\text{ }\|\text{ }\mathrm{MLP}_{x}(\mathbf{x}^{f}_{i}\text{ }\|\text{ }\mathbf{x}^{f,d}_{i})),$ (15)
@@
where $\widetilde{x^{f}_{i}}\in\mathbb{R}^{\Gamma^{\prime}\times|\mathcal{E}|}$, is also optimized by minimizing the corresponding cross-entropy loss. Subsequently, the top-$1$ prediction is generated using the Route Generation Module for the $n=1$ case. The generated future route can be inserted into the first position, replacing the original $K$-th prediction, to obtain the refined top-$K$ future route predictions $\{\widehat{x^{f}_{i,k}}\}_{k=1}^{K}$.
IV-F Multi-Objectives Optimization
@@
where $w_{\cdot}$ are weights for different loss. The complete learning procedure of RouteKG are detailed in Algorithm 2.
@@
1 Randomly initialize $\Theta_{kg}=\{\mathbf{W}_{\mathcal{E}},\mathrm{MLP}_{d},$ $\mathbf{W}_{\mathcal{R}^{c/s/a/d}},$$\mathbf{P}_{\mathcal{R}^{c/s/a/d}}\}\quad\text{and}\quad\Theta_{r}=\{\mathrm{MLP}_{r/f/k/x}\}.$
@@
7       $\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}\leftarrow\mathcal{M}_{g}(\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathbf{G},\mathbf{A})$
8       $\{\widehat{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}\leftarrow\mathcal{M}_{r}(\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K},\mathcal{R};\Theta_{r})$
@@ V-A Data
The Shanghai dataset consists of trajectory records from 10,609 taxis from April 16, 2015, to April 21, 2015, with an average sampling rate of approximately 10 seconds per record. We concentrated on a specific region in Shanghai, adhering to the parameters outlined by [77]. The chosen region’s road network incorporates 320 nodes and 714 links. Each data entry includes the taxi ID, date, time, longitude, latitude, and an occupied flag indicator.
For data preprocessing, we initially employed a fast map-matching algorithm [72] to convert GPS traces into routes on the respective road network. We then cleaned the data, eliminating routes that contained loops and those with too few links (i.e., less than ten links). Subsequently, the refined Chengdu dataset contained 93,125 routes, while the Shanghai dataset comprised 24,468 routes. The road networks of Chengdu and Shanghai are visually represented in Figure 6, and the key network statistics are summarized in Table II.
Figure 6: Selected road networks of Chengdu and Shanghai.
@@
TABLE II: Summary statistics of Chengdu and Shanghai road networks (ID: in-degree, OD: out-degree).
V-B Baseline Methods
In this study, we compare our approach with several established baselines to evaluate performance:
@@ V-B Baseline Methods
  • Dijkstra [37]: Dijkstra’s algorithm is a prominent method for finding the shortest paths between nodes in a graph, where the path length is assumed to be the sum of link length to reflect realistic geographic distance. This baseline can only work when the exact goal location is given.
  • RNN [8]: The RNN is an artificial neural network that recognizes patterns in sequential data. It accomplishes this by utilizing internal memory to process arbitrary sequences of inputs, making it effective for predicting future routes.
  • GRU [10]: The GRU is a type of RNN that utilizes gating mechanisms to capture long-term dependencies in the data, thereby improving the model’s predictive capabilities. This is especially beneficial for applications in trajectory prediction, where long-term dependencies play a crucial role.
  • LSTM [9]: The LSTM is another variant of RNNs that addresses the problem of learning long-term dependencies in data. It does this by implementing a special architecture consisting of a series of memory cells, which effectively control the flow of information, making LSTMs advantageous for predicting future routes.
  • NetTraj [14]: NetTraj is an advanced network-based trajectory prediction model specifically designed for predicting future movements in road networks. It leverages the inherent structure of road networks and utilizes historical trajectory data for accurate predictions. By integrating the Graph Attention Network (GAT) with LSTM, NetTraj offers a robust framework for making future trajectory predictions.
  • RCM-BC [77]: The Route Choice Model-Behavioral Cloning (RCM-BC) is a behavioral cloning approach designed for route choice modeling in sequential decision-making scenarios. It employs supervised learning to create a policy that maps states to actions based on observed behavior to predict future routes. This baseline also requires the knowledge of the exact goal location to work.
@@
TABLE III: Performance comparison of different methods.
V-C Main Results
V-C1 Experimental Settings
To comprehensively assess model performance, we design experiments based on the three subproblems as defined in Section III-A: (1) route prediction with unknown goal $\mathcal{F}_{1}$, (2) route prediction with goal direction only $\mathcal{F}_{2}$, and (3) route prediction with complete goal information $\mathcal{F}_{3}$. We refer to these three subproblems as NoGoal, GoalD, and Goal. They reflect varying degrees of information availability regarding the road user’s intended destination, and represent a broad range of real-world application scenarios. For instance, a system might not know the user’s exact destination due to privacy concerns but could have access to more general information, such as the goal direction.
@@ V-C1 Experimental Settings
Figure 7: Example results in Chengdu (upper) and Shanghai (lower), with the red line indicating the last observed link, the green line the actual future route, and the blue lines are the predicted future routes.
To define evaluation metrics, we consider the top-$k$ predictions for the $i$-th observed route $\{\widehat{x^{f}_{i,k}}\}_{k=1}^{K}$=$\{\{\widehat{e_{i,k}^{j}}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\}_{k=1}^{K}$ and the actual $i$-th future route $x_{i}^{f}=\{e_{i}^{j}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$. The link-level recall R@K is defined as $R@K=\frac{1}{|\mathcal{X}|}\sum_{i=1}^{|\mathcal{X}|}\max_{k=1}^{K}\left[\frac{1}{\Gamma^{\prime}}\sum_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\mathbb{I}(\widehat{e_{i,k}^{j}}=e_{i}^{j})\right].$ where $\mathbb{I}(a=b)=\begin{cases}1&\text{if }a=b\\ 0&\text{otherwise}\end{cases}$ is the indicator function.
Similarly, the route-level recall R@K is defined as $R@K=\frac{1}{|\mathcal{X}|}\sum_{i=1}^{|\mathcal{X}|}\max_{k=1}^{K}\left[\mathbb{I}(\sum_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\mathbb{I}(\widehat{e_{i,k}^{j}}=e_{i}^{j}),\Gamma^{\prime})\right].$
We also compute the route-level MRR of the top-$k$ predictions, M@K, as follows: $M@K=\frac{1}{|\mathcal{X}|}\sum_{i=1}^{|\mathcal{X}|}\sum_{k=1}^{K}\frac{1}{k}\left[\mathbb{I}(\sum_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\mathbb{I}(\widehat{e_{i,k}^{j}}=e_{i}^{j}),\Gamma^{\prime})\right].$
The experiments are conducted on a Ubuntu server with the Python 3.6 environment. The deep learning computations are performed using the PyTorch framework. The server’s hardware specifications include an Intel(R) Xeon(R) Platinum 8375C CPU with a clock speed of 2.90GHz, coupled with 8 NVIDIA GeForce RTX 3090 GPUs, each featuring 24GB of memory. To ensure the robustness and generalizability of our model, hyperparameters are tuned based on the performance of the validation set. The fine-tuning of these hyperparameters is crucial for balancing the bias-variance trade-off and optimizing the model’s performance. All hyperparameters are in VI-B.
V-C2 Main Results Analysis